1. find the probability to the left of z=-2.23 or to the right of z=1.23

2. the number of bases for a baseball player hitter is given here:

number of bases Probability

number of bases=0, probability=.718

number of bases=1 P=.174

bases=2 P=.065

bases=3 P=.004

bases=4 P=.039

2a. find the probability of a hitter getting a double.

2b. find the probability of a hitter getting a double or triple.

8. is this a valid distribution in problem 6? why or why not?

9. find the mean in the table from problem 6.
Answer #1

1)

excel formula for probability from z score is =NORMSDIST(Z)

=P(Z ≤   -2.230   ) =   0.0128737

= P(Z ≥   1.230   ) = P( Z <   -1.230   ) =    0.1093

2)

a)

P (x=2) = 0.065

b)

P (x=2) + P (x=3) = 0.065 +0.004 = 0.069

8)

Distribution is valid as total probability is one.

9)

X P(X) X*P(X)
0 0.7180 0.000
1 0.1740 0.174
2 0.0650 0.130
3 0.0040 0.012
4 0.0390 0.156
P(X) X*P(X)
total sum = 1 0.472

mean = E[X] = Σx*P(X) =            0.4720

Please revert back in case of any doubt.

Please upvote. Thanks in advance.